A 7.5 cm long gold wire has a 0.10 mm cross section diameter. Opposite ends of this
wire are connected to the terminals of a 1.5 V battery.
a) Show the complete calculation of the resistance of the wire.
b) Show the complete calculation of current through the wire.
c) Show the complete calculation of the power supplied by the battery.
1
Expert's answer
2020-02-10T09:42:36-0500

a) "R=\rho*l*4/(\pi*D^2)=0.21 Ohm"

b) "I=U\/R=7.14 A"

c) "P= U*I=10.71 W"